# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 3. Preliminaries on combinatorics
can be written as a $a_{1}b_{1}a_{2}b_{2}a_{3}$, where $a_{1},a_{2},a_{3}\in\left(\begin{matrix}1&p^{m}\mathbb{Z}/p^{n}\mathbb{Z}\\ 0&1\end{matrix}\right)$ and $b_{1},b_{2}\in\left(\begin{matrix}1&0\\ p^{m}\mathbb{Z}/p^{n}\mathbb{Z}&1\end{matrix}\right)$. From this it follows that if we let $q_{1}^{\prime}=\text{gcd}((Q_{1}^{*})^{2},q_{1})$, $q_{2}^{\prime}=\text{gcd}((Q_{2}^{*})^{2},q_{2})$, then $q_{1}^{\prime}q_{2}^{\prime}\leq q^{40\delta}$, and
@@
$A^{720}\supset\Gamma(q^{\prime})/\Gamma(q).$
@@
$A^{7204}\supset\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\ltimes(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z})^{2}.$
@@
Since $|A|>(q_{1}^{3}q_{2}^{2})^{1-\delta}$ and the number of equivalence classes is bounded by $q_{2}^{2}$, there is one class $B$ such that $|B|>q_{1}^{3(1-\delta)}q_{2}^{-2\delta}>q_{1}^{3-5\delta}$. Take any $b\in B$, then $Bb^{-1}>q_{1}^{3-5\delta}$ and the unipotent part of $Bb^{-1}$ is 00. Applying Lemma 3.19, we have
$A^{1440}\supset(Bb^{-1})^{720}\supset\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$ (3.21)
@@
By the pigeon hole principle, there exists $\gamma_{0}\in\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})$ such that $B^{\prime}=\{(\gamma_{0},w),w\in\mathbb{Z}/q_{2}\mathbb{Z}\}\cap A$ has at least $\frac{|A|}{q_{1}^{3}}=q_{1}^{-3\delta}q_{2}^{2-2\delta}$. Take any $b^{\prime}\in B^{\prime}$. It follows that there exists $(1,w)\in B^{\prime}b^{\prime-1}$, $w=q^{\prime}w_{0}$, where $q^{\prime}|q_{2},q^{\prime}\leq q_{1}^{3\delta}q_{2}^{2\delta}$ and $w_{0}$ is a primitive vector in $(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$.
@@
Let $L\subset A^{2882}$ be the collection of conjugates of $(1,w_{0})$ by $\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$. Using (3.22), we obtain
$A^{5764}\supset LL^{-1}\supset(1,\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})\mathbb{Z}/q_{2}\mathbb{Z})$ (3.23)
@@
Let $u,v\in\mathbb{Z}$ and $f_{u,v}:\mathbb{Z}^{6}\rightarrow\mathbb{Z}^{2}$ be given by
@@
It is an elementary exercise to check
$H_{u,v}:=\left\{\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),\left(\begin{matrix}E\\ F\end{matrix}\right)\right)\in G\bigg{|}f_{u,v}=0\right\}.$
is a subgroup of $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$.
The following lemma shows if a sectional map from (a large subgroup of) $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$ to $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}$ is a homomorphism, then the image must be close to $H_{u,v}$.
@@
Let $1\leq m<\frac{n}{4}$ and let $H$ be a subgroup of $\Gamma(p^{m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}$ such that $\mathbb{P}_{0}(H)=\Gamma(p^{m})/\Gamma(p^{n})$ and $|H|=|\Gamma(p^{m})/\Gamma(p^{n})|$. Then there exist some $u,v\in\frac{1}{p^{m}}\mathbb{Z}$ such that
$H\cap\left(\Gamma(p^{2m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}\right)(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor})\subset H_{u,v}(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor})$
@@
Lemma 3.24 follows from proofs of a series of claims.
@@
$(1+g+g^{2}+\cdots+g^{p-1})w\equiv(g^{p},p(1+\frac{(p-1)}{2}p^{k}M)w)(\operatorname{mod}{p^{\min\{2k,n-m\}}}).$
@@
$(1+\frac{p-1}{2}p^{k}M)pw\in(p^{k-m+1}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}.$
@@
Claim 3: If $41\leq m\leq n/$ 4, $\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ v_{1}\end{matrix}\right)\right)\in H$, then $v_{1}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}$.
@@
If taking $t=\lfloor\frac{n}{4}\rfloor-m$, by Claim 2 we get $v_{1}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}.$ This finishes the proof of Claim 3.
@@
Claim 4: If $41\leq m\leq n/$ 4, $\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}u_{2}\\ v_{2}\end{matrix}\right)\right)\in H$, then $u_{2}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}$.
Proposition 3.24 thus follows by taking $u=\frac{1}{p^{m}}v_{2}$ and $v=\frac{1}{p^{m}}u_{1}$ since
$\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ \cdot\end{matrix}\right)\right),\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}\cdot\\ v_{2}\end{matrix}\right)\right)\in H_{u,v}(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor}),$

### v4

@@ 3. Preliminaries on combinatorics
can be written as a $a_{1}b_{1}a_{2}b_{2}a_{3}$, where $a_{1},a_{2},a_{3}\in\left(\begin{matrix}1&p^{m}\mathbb{Z}/p^{n}\mathbb{Z}\\ 0&1\end{matrix}\right)$ and $b_{1},b_{2}\in\left(\begin{matrix}1&0\\ p^{m}\mathbb{Z}/p^{n}\mathbb{Z}&1\end{matrix}\right)$. From this it follows that if we let $q_{1}^{\prime}=\text{gcd}((Q_{1}^{\prime})^{2},q_{1})$, $q_{2}^{\prime}=\text{gcd}((Q_{2}^{\prime})^{2},q_{2})$, then $q_{1}^{\prime}q_{2}^{\prime}\leq q^{40\delta}$, and
@@
$A^{2880}\supset\Gamma(q^{\prime})/\Gamma(q).$
@@
$A^{28804}\supset\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\ltimes(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z})^{2}.$
@@
Since $|A|>(q_{1}^{3}q_{2}^{2})^{1-\delta}$ and the number of equivalence classes is $q_{2}^{2}$, there is one class $B$ such that $|B|>q_{1}^{3(1-\delta)}q_{2}^{-2\delta}>q_{1}^{3-5\delta}$. Take any $b\in B$, then $Bb^{-1}>q_{1}^{3-5\delta}$ and the unipotent part of $Bb^{-1}$ is 0. Applying Lemma 3.19, we have
$A^{5760}\supset(Bb^{-1})^{2880}\supset\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$ (3.21)
@@
By the pigeon hole principle, there exists $\gamma_{0}\in\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})$ such that $B^{\prime}=\{(\gamma_{0},w),w\in\mathbb{Z}/q_{2}\mathbb{Z}\}\cap A$ has at least $\frac{|A|}{q_{1}^{3}}=q_{1}^{-3\delta}q_{2}^{2-2\delta}$ elements. Take any $b^{\prime}\in B^{\prime}$. It follows that there exists $(1,w)\in B^{\prime}b^{\prime-1}$, $w=q^{\prime}w_{0}$, where $q^{\prime}|q_{2},q^{\prime}\leq q_{1}^{3\delta}q_{2}^{2\delta}$ and $w_{0}$ is a primitive vector in $(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$.
@@
Let $R\subset A^{11522}$ be the collection of conjugates of $(1,w_{0})$ by $\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$. Using (3.22), we obtain
$A^{23044}\supset RR^{-1}\supset(1,\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})\mathbb{Z}/q_{2}\mathbb{Z})$ (3.23)
@@
Let $u,v\in\mathbb{Q}$ and $f_{u,v}:\mathbb{Z}^{6}\rightarrow\mathbb{Z}^{2}$ be given by
@@
It is straightforward that
$H_{u,v}:=\left\{\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),\left(\begin{matrix}E\\ F\end{matrix}\right)\right)\in G\bigg|f_{u,v}(A,B,C,D,E,F)=0\right\}.$
is a subgroup of $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ as $H_{u,v}$ is the stabilizer of $(u,v)^{t}$.
The following proposition shows if a sectional map from (a large subgroup of) $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$ to $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}$ is a homomorphism, then the image must be close to $H_{u,v}$ for some $u,v$.
@@
Let $1\leq m<\frac{n}{4}$ and let $H$ be a subgroup of $\Gamma(p^{m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}$ such that $\mathbb{P}_{0}(H)=\Gamma(p^{m})/\Gamma(p^{n})$ and $|H|=|\mathbb{P}_{0}(H)|$. Then there exist some $u,v\in\frac{1}{p^{m}}\mathbb{Z}$ such that
$H\cap\left(\Gamma(p^{2m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}\right)(\operatorname{mod}p^{[\frac{n}{4}]})\subset H_{u,v}(\operatorname{mod}p^{[\frac{n}{4}]})$
@@
Proposition 3.24 follows from proofs of a series of claims.
@@
$(1+g+g^{2}+\cdots+g^{p-1})w\equiv p(1+\frac{(p-1)}{2}p^{k}M)w(\operatorname{mod}{p^{\min\{2k,n-m\}}}).$
@@
$p(1+\frac{p-1}{2}p^{k}M)w\in(p^{k-m+1}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}.$
@@
Claim 3: If $1\leq m\leq n/4$, $\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ v_{1}\end{matrix}\right)\right)\in H$, then $v_{1}\equiv 0\pmod{p^{[\frac{n}{4}]}}$.
@@
If taking $t=[\frac{n}{4}]-m$, by Claim 2 we get $v_{1}\equiv 0\pmod{p^{[\frac{n}{4}]}}$. This finishes the proof of Claim 3.
@@
Claim 4: If $1\leq m\leq n/4$, $\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}u_{2}\\ v_{2}\end{matrix}\right)\right)\in H$, then $u_{2}\equiv 0\pmod{p^{[\frac{n}{4}]}}$.
Proposition 3.24 thus follows by taking $u=\frac{1}{p^{m}}v_{2}$ and $v=\frac{1}{p^{m}}u_{1}$, since
$\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ \cdot\end{matrix}\right)\right),\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}\cdot\\ v_{2}\end{matrix}\right)\right)\in H_{u,v}(\operatorname{mod}p^{[\frac{n}{4}]}),$
@@
Finally, we record some elementary combinatorics on $\text{SL}_{2}$.
Lemma 3.25.
Let $H_{1},H_{2}\subset\text{SL}_{2}(\mathbb{Z})$ and let $p$ be a prime. Suppose $1+p^{m_{1}}V(\operatorname{mod}p^{m_{2}})\subset H_{1}(\operatorname{mod}p^{m_{2}})$ and $1+p^{n_{1}}V(\operatorname{mod}p^{n_{2}})\subset H_{2}(\operatorname{mod}p^{n_{2}})$ with $1\leq m_{1}\leq m_{2}\leq 2m_{1}$ and $1\leq n_{1}\leq n_{2}\leq 2n_{1}$. Then
$1+2p^{m_{1}+n_{1}}V\subset[H_{1},H_{2}]^{3}(\operatorname{mod}p^{\min\{m_{2}+n_{2}-1,2m_{1}+n_{1},m_{1}+2n_{1}\}}).$ (3.26)
To prove Lemma 3.25, we first need
Lemma 3.27.
Let $p$ be a prime. Let $A,B\in\mathbb{Z},m_{1},m_{2}\in\mathbb{N}$ such that
$\pi_{p^{m_{1}}}(A)\supset\mathbb{Z}/p^{m_{1}}\mathbb{Z}$
$\pi_{p^{m_{2}}}(B)\supset\mathbb{Z}/p^{m_{2}}\mathbb{Z}.$
Then
$\pi_{p^{m_{1}+m_{2}-1}}\left(\sum_{2}AB\right)\supset\mathbb{Z}/p^{m_{1}+m_{2}-1}\mathbb{Z}.$
Proof of Lemma 3.27.
Take $a\in A$ such that $v_{p}(a)=m_{1}-1$, then
$aB(\operatorname{mod}p^{m_{1}+m_{2}-1}\mathbb{Z})=p^{m_{1}-1}\mathbb{Z}/p^{m_{1}+m_{2}-1}\mathbb{Z}.$ (3.28)
Take $b\in B$ such that $v_{p}(b)=0$, then
$Ab(\operatorname{mod}p^{m_{1}-1}\mathbb{Z})=\mathbb{Z}/p^{m_{1}-1}\mathbb{Z}.$ (3.29)
(3.28) and (3.29) then implies $aB+Bb(\operatorname{mod}p^{m_{1}+m_{2}-1})=\mathbb{Z}/p^{m_{1}+m_{2}-1}\mathbb{Z}$. ∎
Proof of Lemma 3.25.
Take $g_{1}\in H_{1}$ and $g_{2}\in H_{2}$. Write
$g_{1}=1+p^{m_{1}}\begin{pmatrix}a_{1}&b_{1}\\ c_{1}&-a_{1}\end{pmatrix}+p^{2m_{1}}M_{1},$
$g_{2}=1+p^{m_{2}}\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&-a_{2}\end{pmatrix}+p^{2m_{2}}M_{2},$
with $a_{1},b_{1},c_{1}\in[0,p^{m_{1}}-1]$, $a_{2},b_{2},c_{2}\in[0,p^{m_{2}}-1]$, and $M_{1},M_{2}\in\text{Mat}_{2}(\mathbb{Z})$.
By Lemma 5.66, we get
$g_{1}g_{2}g_{1}^{-1}g_{2}^{-1}$ $\equiv 1+[g_{1},g_{2}]\equiv 1+p^{m_{1}+n_{1}}\bigg[\begin{pmatrix}a_{1}&b_{1}\\ c_{1}&-a_{1}\end{pmatrix},\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&-a_{2}\end{pmatrix}\bigg$]
$=1+p^{m_{1}+n_{1}}\begin{pmatrix}b_{1}c_{2}-b_{2}c_{1}&2(a_{1}b_{2}-b_{1}a_{2})\\ 2(c_{1}a_{2}-a_{1}c_{2})&-(b_{1}c_{2}-b_{2}c_{1})\\ \end{pmatrix}\pmod{p^{m_{1}+n_{1}+\min\{m_{1},n_{2}\}}}$.
If we take $a_{1},a_{2}=0$ and $b_{1},c_{1},b_{2},c_{2}$ arbitrary, then we can apply Lemma 3.27 to obtain
$[H_{1},H_{2}]\supset 1+p^{m_{1}+n_{1}}\mathbb{Z}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\pmod{p^{\min\{m_{2}+n_{2}-1,2m_{1}+n_{1},m_{1}+2n_{1}\}}}$. (3.30)
Similarly, by setting $b_{1}=0$ and $b_{2}=0$, we can get
$[H_{1},H_{2}]$ $\supset 1+2p^{m_{1}+n_{1}}\mathbb{Z}\begin{pmatrix}0&0\\ 1&0\end{pmatrix}\pmod{p^{\min\{m_{2}+n_{2}-1,2m_{1}+n_{1},m_{1}+2n_{1}\}}}$, (3.31)
and by setting $c_{1},c_{2}=0$, we get
$[H_{1},H_{2}]$ $\supset 1+2p^{m_{1}+n_{1}}\mathbb{Z}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}\pmod{p^{\min\{m_{2}+n_{2}-1,2m_{1}+n_{1},m_{1}+2n_{1}\}}})$ (3.32)
Then (3.26) follows from (3.30), (3.31) and (3.32). ∎
Lemma 3.33.
Let $p$ be a prime, $a,b\in\mathbb{N}$ and $b\geq 2a\geq 0$. Suppose $H\subset\text{SL}_{2}(\mathbb{Z})$ and $\Gamma(p^{a})(\operatorname{mod}p^{b})=H(\operatorname{mod}p^{b})$.
  1. (1)
    If $a\neq 0$, then if $p\neq 2$,
    $[H,H]^{O(\log\frac{b}{a})}(\operatorname{mod}p^{b})=\Gamma(p^{2a})/\Gamma(p^{b}).$ (3.34)
    If $p=2$,
    $[H,H]^{O(\log\frac{b}{a})}(\operatorname{mod}2^{b})=\Gamma(2^{2a+1})/\Gamma(2^{b}).$ (3.35)
  2. (2)
    If $a=0$, then
    $[H,H]^{O(\log b)}(\operatorname{mod}p^{b})=\Gamma_{p^{b}}.$ (3.36)
Proof.
Assume $p\neq 2$.
If $a\neq 0$, let $s_{0}$ be the largest integer such that $1.5^{s_{0}}<\frac{b}{a}$. Apply Lemma 3.25 to $H_{1}=H_{2}=H$ with $m_{1}=n_{1}=[a\cdot 1.5^{j}]$, $m_{2}=n_{2}=\min\{2m_{1},b\}$, $0\leq j\leq s_{0}$. Multiplying the implied sets together yields (3.34).
If $a=0$, Then we have
$[H,H]^{3}(\operatorname{mod}p)=\Gamma_{p}.$ (3.37)
See Theorem 1.1 of [Sha09]. Let $H_{0}\subset[H,H]^{3}$ be the implied set of representatives of $\Gamma_{p}$. It follows from Lemma 4 of [BV12] that $H_{0}(\operatorname{mod}p^{2})$ is not a group, so there is $h_{1},h_{2},h_{3}\in H_{0}$ such that $h_{1}h_{2}h_{3}^{-1}\equiv 1(\operatorname{mod}p)$ but $h_{1}h_{2}h_{3}^{-1}\not\equiv 1(\operatorname{mod}p^{2})$. From here it is easy to deduce that $H_{0},H_{0}]^{O(1)}(\operatorname{mod}p^{2})=\Gamma_{p^{2}}$.
We also know
$[H,H]^{O(\log b)}(\operatorname{mod}p^{b})\supset\Gamma_{p^{b}}$ (3.38)
by the $a=1$ case of (3.34). Combining (3.37) and (3.38) gives (3.36).
The proof for the $p=2$ case is analogous.
Lemma 3.39.
Let $p$ be a prime and $b\geq 2a\geq 1$. Suppose $\gamma_{0}\in\text{SL}_{2}(\mathbb{Z})$ with $v_{p}(\gamma_{0}-1)=a$, and $H\subset SL_{2}(\mathbb{Z})$ with $H\supset 1+p^{b}V(\operatorname{mod}p^{2b})$. Then
$\gamma_{0},H],H]^{2}\supset 1+4p^{a+2b}V(\operatorname{mod}p^{a+3b}).$
Proof.
Assume $p$ is odd. Write $\gamma_{0}\equiv 1+p^{a}X_{0}(\operatorname{mod}p^{2a})$ for some traceless $X_{0}\in\text{Mat}_{2}(\mathbb{Z})$ that is also primitive mod $p$.
It is an elementary exercise to check that it is always possible to choose $h_{1},h_{2}\in H$, $h_{1}\equiv 1+p^{b}Y_{1}(\operatorname{mod}p^{2b}),h_{2}\equiv 1+p^{b}Y_{2}(\operatorname{mod}p^{2b})$, $Y_{1},Y_{2}$ traceless and primitive mod $p$, so that $X_{1}=X_{0}Y_{1}-Y_{1}X_{0}$ and $X_{2}=X_{0}Y_{2}-Y_{2}X_{0}$ are primitive and are not colinear mod $p$.
Write $\gamma_{1}=\gamma_{0}h_{1}\gamma_{0}^{-1}h_{1}^{-1}$ and $\gamma_{2}=\gamma_{0}h_{2}\gamma_{0}^{-1}h_{2}^{-1}$. From Lemma 5.66,
$\gamma_{1}\equiv 1+\gamma_{0}h_{1}-h_{1}\gamma_{0}\equiv 1+p^{a+b}X_{1}(\operatorname{mod}p^{2a+b})$ (3.40)
$\gamma_{2}\equiv 1+\gamma_{0}h_{2}-h_{2}\gamma_{0}\equiv 1+p^{a+b}X_{2}(\operatorname{mod}p^{2a+b})$ (3.41)
Take $h_{3},h_{4}\in H$ and write $h_{3}\equiv 1+p^{b}Y_{3}(\operatorname{mod}p^{2b})$ and $h_{4}\equiv 1+p^{b}Y_{4}(\operatorname{mod}p^{2b})$.
Applying Lemma 5.66 again,
$\gamma_{1}h_{3}\gamma_{1}^{-1}h_{3}^{-1}\cdot\gamma_{2}h_{4}\gamma_{2}^{-1}h_{4}^{-1}\equiv 1+p^{a+2b}([X_{1},Y_{3}]+[X_{2},Y_{4}])(\operatorname{mod}p^{a+3b}).$ (3.42)
Since $X_{1},X_{2}$ are primitive and not colinear mod $p$, it is another elementary exercise to check that
$\{[X_{1},Y_{3}]+[X_{2},Y_{4}]:Y_{3},Y_{4}\in V\}(\operatorname{mod}p^{b})=V(\operatorname{mod}p^{b}),$ (3.43)
(Hint: it suffices to assume $X_{1}=\left(\begin{matrix}0&1\\ 0&0\end{matrix}\right)$ and $X_{2}=\left(\begin{matrix}0&0\\ 1&0\end{matrix}\right)$, by identifying $V$ with $\mathbb{Z}^{3}$ and use a matrix in $\text{SL}_{3}(\mathbb{Z})$ to change variables.) (3.42) and (3.43) then imply Lemma 3.39.
The $p=2$ case follows from the same argument. The extra factor 4 is due to the fact that $2V\subset[V,V]\neq V$.
Lemma 3.44.
Let $p_{1},p_{2}$ be two primes. Suppose $f$ is a homomorphism from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ for some $0\leq m_{1}\leq n_{1},1\leq m_{2}\leq n_{2}$, and for some $\xi\in\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$, $f(\xi)\neq 1$. Then $p_{1}=p_{2}$ and $v_{p_{1}}(\xi-1)\leq m_{1}+n_{2}-m_{2}$.
Proof.
If $m_{1}=0$, and let us suppose $p_{1}\neq p_{2}$, then $f$ will factor through $\Gamma(p_{1})/\Gamma(p_{1}^{n_{1}})$, which induces a homomorphism $f^{\prime}:\Gamma_{p_{1}}\rightarrow\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$. Since $\Gamma_{p_{1}}$ is almost simple, $f^{\prime}$ must either be trivial, or injective, or factor through the center $Z$. $f^{\prime}$ can not be injective or factor through the center because $\left(\begin{matrix}1&1\\ 0&1\end{matrix}\right)\in\Gamma_{p_{1}}$ has order $p_{1}$, which forces $p_{1}|p_{2}$ if $f$ is injective or factor through $Z$, a contradiction. Hence, $f^{\prime}$ as well as $f$ must be trivial. Therefore, if $f$ is nontrivial, we must have $p_{1}=p_{2}$.
Suppose $p_{1}=p_{2}=p$ and $f(\xi)=y\not\equiv 1(\operatorname{mod}p^{n_{2}})$. Since the order of all elements in $\Gamma(p^{m_{2}})/\Gamma(p^{n_{2}})$ is bounded by $p^{n_{2}-m_{2}}$, it follows that $\xi$ can not be a $p^{n_{2}-m_{2}}$-power in $\Gamma(p^{m_{1}})/\Gamma(p^{n_{1}})$, and this will force $v_{p}(\xi)\leq m_{1}+n_{2}-m_{2}$. ∎